# SuperGrasp: Single-View Object Grasping via Superquadric Similarity Matching, Evaluation, and Refinement

### 31 Mar 2026 v1

@@ Abstract
Robotic grasping from single-view observations remains a critical challenge in manipulation. Existing methods still struggle to generate stable and valid grasp poses when confronted with incomplete geometric information. To address these limitations, we propose SuperGrasp, a novel two-stage framework for single-view grasping with parallel-jaw grippers that decomposes the grasping process into initial grasp pose generation and subsequent grasp evaluation and refinement. In the first stage, we introduce a Similarity Matching Module that efficiently retrieves grasp candidates by matching the input single-view point cloud with a pre-computed primitive dataset based on superquadric coefficients. In the second stage, we propose E-RNet, an end-to-end network that expands the grasp-aware region and takes the initial grasp closure region as a local anchor region, enabling more accurate and reliable evaluation and refinement of grasp candidates. To enhance generalization, we construct a primitive dataset containing 1.5k primitives for similarity matching and collect a large-scale point cloud dataset with 100k stable grasp labels from 124 objects for network training. Extensive experiments in both simulation and real-world environments demonstrate that our method achieves stable grasping performance and strong generalization across varying scenes and novel objects.
@@ I Introduction
Grasping is a fundamental capability in robotic manipulation and has shown great potential in household service and industrial production-line applications in structured environments. However, grasping objects in unstructured environments remains challenging, especially under single-view observations, due to incomplete and noisy visual inputs. Effectively exploiting such degraded geometric information to generate stable grasp poses is still difficult for existing methods.
Superquadrics[1] were introduced into computer vision in the late twentieth century as a compact representation of geometric primitives. Later, with improvements to fitting algorithms by[13, 22], superquadrics became increasingly effective for representing object geometry directly from point clouds. This progress has inspired new solutions for robotic grasping. For example, Wu et al.[29] take point clouds as input, compute multiple graspable regions based on superquadric coefficients, and then sample grasp poses accordingly, achieving promising results. However, these methods are generally developed under the assumption of complete point clouds. For single-view point clouds, geometric representation still suffers from substantial difficulties in both shape and scale alignment.
With the rapid development of neural networks, data-driven grasping methods, including sampling-and-evaluation pipelines[25, 12, 30] and end-to-end approaches[18, 24, 30, 5, 31, 27, 11], have achieved significant progress in this field. Representative methods such as GPD[25] and PointNetGPD[12] adopt a two-stage framework. They first sample grasp candidates from point-cloud geometry and then classify these candidates using neural networks. However, due to geometric incompleteness in sparse point clouds, the sampled grasp candidates are often unreliable, and the subsequent evaluation models may lack robustness. To address this issue, end-to-end methods[24, 27, 26, 14, 5, 31, 20] directly predict grasp poses from input point clouds. Although effective, these methods usually require large-scale training data to learn feasible grasp poses, and their generalization to unseen objects is still limited.
@@
Inspired by chen et al.[3], we propose SuperGrasp, a novel two-stage grasping pipeline, to address the above challenges through a similarity-matching module and a more robust evaluation and refinement module. Specifically, in the first stage, we estimate superquadric coefficients from the target object point cloud under a single-view observation to represent its geometric shape. We then perform similarity matching against a primitive database constructed offline using superquadric coefficients, where both shape and scale are considered. The grasp poses of the top-$N$ most similar primitive objects are subsequently transformed to the target object point cloud, yielding rich and effective grasp candidates.
In the second stage, to enable more reliable evaluation and refinement, we design a new evaluation-refinement network based on PointNet++[23] as the backbone. The network enlarges the grasp-aware region and takes the initial grasp closure local region as an anchor region. Based on this design, it jointly analyzes the geometric features within the initial grasp closure region and the surrounding spatial geometric context, thereby enabling more accurate and robust prediction of grasp feasibility as well as refinement feasibility.
To validate the effectiveness of the proposed method, we conduct extensive grasping experiments in both simulation and real-world environments across multiple scenarios. The experimental results demonstrate that our method achieves strong stability while maintaining excellent generalization capability. The task pipeline is shown in Fig. 1. Our main contributions are summarized as follows:
  • We propose a superquadric-based similarity matching method. It represents a target object with incomplete point clouds using multiple primitive objects with similar shapes and scales from a complete point cloud database and transfers high-quality grasp candidates from the matched primitives to the target object, thereby providing rich and effective grasp candidates.
  • We propose a new evaluation-refinement network that expands the grasp-aware region and takes the initial grasp closure local region as an anchor region. This design enables the network to analyze not only the geometric features within the grasp closure region but also the geometric relationships of the surrounding points, leading to more robust prediction of grasp feasibility and refinement feasibility.
@@ II Related Work
Grasping has long been a focus of robotics research. With the rapid development of computer vision and neural networks, grasping methods have become increasingly diverse. In this section, we introduce superquadrics and discuss two primary research directions in robotic grasping.
@@ II-A Superquadrics
To fully characterize a superquadric in 3D space, 5 shape parameters and a 6-DoF pose are required:
$\{a_{x},a_{y},a_{z},\epsilon_{1},\epsilon_{2},\bm{g}\},$
where $\bm{g}=(\bm{R},\bm{t})$ denotes the pose of the superquadric, R{R} represents its orientation, and t{t} denotes its translation. Superquadrics are widely applied in object modeling, shape approximation, and scene understanding due to their low-dimensional representation and strong geometric interpretability. They are particularly effective for modeling objects with nearly convex structures, and the recovered parameters can directly convey the graphical, orientational, and shape characteristics of the objects.
@@ II-B Two-Stage Grasping
Two-stage grasping methods typically decompose the problem into grasp candidate generation and grasp evaluation. GPD[25] first samples grasp candidates in 3D space, labels them using force closure analysis [19], and trains a CNN to predict grasp feasibility from grasp-aligned projections. PointNetGPD[12] further improves this paradigm by directly encoding the point cloud within the grasp region using a PointNet-based architecture for grasp quality prediction. Later, 3DSGrasp[16] introduced 3D reconstruction to compensate for missing geometric information in point clouds, thus improving both the generation of grasp candidates and the accuracy of grasp feasibility analysis. However, due to the continuous and high-dimensional nature of the grasp space, exhaustive spatial sampling is inherently time-consuming and inefficient.
To address this issue, Chen et al.[3] proposed to generate grasp candidates through multi-Level similarity matching. By measuring the geometric and semantic similarity between the target object and objects stored in a database, grasp poses from the most similar objects can be transferred to the target object. Nevertheless, under incomplete geometric observations, the matching quality of such methods often degrades significantly.
@@ II-C End-to-End Grasping
To address sampling inefficiency, Fang et al. [5] proposed an end-to-end network frameworks that decouple and discretize grasp representations. These models take point clouds as input and directly output grasp candidates with quality scores, relying on large-scale training data to ensure strong generalization. Methods [27, 31] further introduced grasp refinement strategies, first analyzing global point cloud information to obtain coarse grasp poses, then refining them using local geometric details, fully leveraging point cloud spatial information and improving grasp success rates. Methods [17, 6] proposed VAE-based models that encodes point clouds into implicit features and decodes them to generate 6-DoF grasp poses. However, the aforementioned methods still exhibit limited generalization capabilities when handling variations in scene and object shapes.
@@ III Problem Formulation
We formulate the task as tabletop decluttering with a robotic arm equipped with a parallel-jaw gripper, although the proposed framework can also be applied to target-object grasping. Notably, only a single fixed depth camera is deployed within a $50\,\mathrm{cm}\times 50\,\mathrm{cm}$ workspace to capture the scene point cloud. The model takes as input the target object point cloud together with the global scene point cloud $X\in\mathbb{R}^{N\times 3}$, and outputs a grasp pose $g$, defined as:
@@ IV-A Overview
Single-view grasping remains challenging due to missing geometric information and sensor noise. To improve the efficiency and reliability of grasping, we propose SuperGrasp, a new two-stage framework. First, we introduce a geometric similarity matching strategy to generate grasp candidates(see Sec. IV-B). Second, we propose an end-to-end evaluation and refinement network, termed E-RNet, to evaluate the stability of the grasp and refine it locally(see Sec. IV-D).
Figure 3: The primitive object database constructed in this work, which contains 1.5k objects of various shapes and scales, including cylinders, frustums, elliptical cylinders, and cuboids.
@@ IV-B Superquadric Similarity Matching
Superquadrics can represent a complete object shape using only six parameters $\{a_{x},a_{y},a_{z},\epsilon_{1},\epsilon_{2},\bm{g}\}$, providing an alternative grasping paradigm. However, in the single-view setting, point clouds are often unclosed, making it difficult to accurately estimate superquadric coefficients. This leads to biases in both shape and scale, resulting in unstable grasp candidate quality and insufficient coverage. To address this issue, we propose a superquadric-based similarity matching method, as illustrated in Fig.4.
@@
We first build a database consisting of primitive objects (e.g., cuboids, cylinders, frustums), as shown in Fig.3. To better fit object scales, we apply proportional scaling to all primitives. For each object, we precompute a set of stable pre-grasp poses $G_{1}^{n}=\{g_{1},g_{2},\ldots\}$ and fix the gripper width $w$, defined as the distance from the inner finger surfaces to the object contact surfaces (set to $0.0125\,\mathrm{m}$ in our experiments). The database contains $1.5\mathrm{k}$ objects in total.
@@
To accelerate retrieval, we precompute superquadric coefficients using the EMS (Expectation-Maximization Search) fitting method[13] for all database objects and store them in a hash table. Given a target object, we likewise use EMS to fit its superquadric and obtain the scale parameters $(a_{x},a_{y},a_{z})$ and shape parameters $(\epsilon_{1},\epsilon_{2})$. To separately quantify overall size and aspect ratio, we calculated the geometric mean of the following proportional parameters as follows:
$g=(a_{x}a_{y}a_{z})^{1/3},$ (2)
@@
$r=\left[\frac{a_{x}}{g},\frac{a_{y}}{g},\frac{a_{z}}{g}\right].$ (3)
@@
Figure 4: Overview: Our framework consists of two stages. In the first stage, we first estimate the superquadric coefficients of the target object and compute their similarity to those of the objects in the database. We then select the top-$N$ most similar objects and transfer their grasp poses to the target object through a transformation matrix $T$, followed by a coarse filtering step to remove noisy candidates. In the second stage, we enlarge the perception region of each initial grasp and feed the cropped regional point cloud into a PointNet++ network for feature extraction. We then further crop the features corresponding to the points within the initial gripper closing region and aggregate them via max pooling to obtain the grasp anchor feature. Finally, the anchor feature is fed into an evaluation network and a refinement network to predict grasp feasibility and refinement feasibility, respectively.
@@
$d_{\mathrm{scale}}=\left|\log\frac{g^{(q)}}{g^{(c)}}\right|,$ (7)
@@
We rank the remaining candidates in ascending order of $s$ and select the top-$K_{2}$ most similar objects. Their grasp candidates are then transferred to the target object frame using $T$, the relative pose between the fitted candidate and query superquadrics. In all experiments, we set $w_{r}=1.0$, $w_{\epsilon}=1.2$, $\lambda_{s}=1.6$, $\lambda_{a}=2.0$, $K_{1}=50$ and $K_{2}=5$.
@@
We perform an additional coarse filtering step to remove unreasonable candidates, including: i) those for which the gripper collides with the target points, ii) those that contain fewer than 50 target points between the gripper fingers, and iii) those for which the angle between the gripper closing direction and the surface normal at the fingertip contact points exceeds $20^{\circ}$.
This module yields rich and reliable grasp candidates for the subsequent evaluation-and-refinement model.
@@ IV-C Data Collection
For label assignment, unlike methods[12, 25], which grade grasp poses based on analytic metrics, we directly determine the label according to the execution outcome in simulation. Specifically, during execution, if the gripper collides with surrounding objects, the label is directly set to 0. If the gripper successfully lifts the target object, then reorients it to a pose with the gripper approach axis aligned vertically downward, and the grasp remains stable under a downward disturbance force of $1\,\mathrm{N}$ applied for $0.5\,\mathrm{s}$, as shown in Fig. 5, the label is set to 1. All other cases are labeled as 0.
@@ IV-D E-RNet: Evaluation and Refinement Network
between the gripper fingers and the local surface normals of the object, a certain degree of relative motion may occur between the gripper and the object during closure. Therefore, it is insufficient to consider only the grasp closing region; the spatial information around this region is also important for reliable grasp evaluation. In addition, although the first stage provides abundant grasp candidates, dense clutter and geometric incompleteness may still cause all initial candidates to fail, in which case slight pose adjustment can play a critical role. To address these issues, we propose a novel evaluation and refinement network, termed E-RNet. The overall architecture is illustrated in Fig. 4.
To enrich grasp-aware perception, we enlarge the initial gripper closing region. Specifically, the initial grasp pose is first deepened by $0.008\,\mathrm{m}$ along the approach direction, and then additionally rotated around the approach axis by $15^{\circ}$ and $15^{\circ}$. For stable learning, we transform the global point cloud into the gripper coordinate frame and use the enlarged closing region to crop the local region of interest from the global point cloud, yielding a cropped point cloud of size $N\times 3$. The cropped point cloud is then fed into a PointNet++ backbone to extract point-wise features of size $N\times 128$.
@@
we use a two-class softmax loss for grasp feasibility classification.
Unlike methods that fine-tune all grasp candidates in a multi-dimensional pose space to further expand the candidate set[10], we instead take the initial grasp region as a spatial anchor and explore the grasp feasibility of its neighboring regions, which is more efficient and time-saving. Moreover, unlike Regnet[31], which regresses each candidate towards a single optimal target, our training strategy allows multiple valid refinement choices, as long as they can lead to a stable grasp. This design provides more refinement flexibility during inference.
@@
$\mathcal{L}_{\mathrm{r}}=$ $-\frac{1}{3N}\sum_{i=1}^{N}\sum_{j=1}^{3}\Big[y_{ij}\log\sigma(z_{ij})$ (11)

### 15 Apr 2026 v2

@@ Abstract
Robotic grasping from single-view observations remains a critical challenge in manipulation. However, existing methods still struggle to generate reliable grasp candidates and stably evaluate grasp feasibility under incomplete geometric information. To address these limitations, we present SuperGrasp, a new two-stage framework for single-view parallel-jaw grasping. In the first stage, we introduce a Similarity Matching Module that efficiently retrieves valid and diverse grasp candidates by matching the input single-view point cloud with a precomputed primitive dataset based on superquadric coefficients. In the second stage, we propose E-RNet, an end-to-end network that expands the grasp-aware region and takes the initial grasp closure region as a local anchor region, capturing the contextual relationship between the local region and its surrounding spatial neighborhood, thereby enabling more accurate and reliable grasp evaluation and introducing small-range local refinement to improve grasp adaptability. To enhance generalization, we construct a primitive dataset containing 1.2k standard geometric primitives for similarity matching and collect a point cloud dataset of 100k samples from 124 objects, annotated with stable grasp labels for network training. Extensive experiments in both simulation and real-world environments demonstrate that our method achieves stable grasping performance and good generalization across novel objects and clutter scenes.
@@ I Introduction
Grasping is a fundamental capability in robotic manipulation and has shown great potential in household service and industrial production-line applications in structured environments. However, grasping objects in unstructured environments remains challenging, especially under single-view observations, due to incomplete and noisy visual input. Using such degraded geometric information to generate stable grasp poses is still difficult for existing methods.
Superquadrics[1] were introduced into computer vision in the late twentieth century as a compact representation of geometric primitives. Later, with improvements to fitting algorithms by[14, 22], superquadrics became increasingly effective for representing object geometry directly from point clouds. This progress has inspired new solutions for robotic grasping. For example, Wu et al.[29] take point clouds as input, compute multiple graspable regions based on superquadric coefficients, and then sample grasp poses accordingly, achieving promising results. However, these methods are generally developed under the assumption of complete point clouds. For single-view point clouds, geometric representation still suffers from substantial difficulties in both shape and scale alignment. To alleviate this issue, Chen et al.[3] proposed a multi-level similarity matching strategy to address the difficulty of grasp candidate generation under single-view observations. Nevertheless, similarity estimation can be dominated by visible local surfaces, while the unseen global geometry is weakly constrained. As a result, the transferred grasp may fit the observed region, but still collide with the invisible part of the target object.
With the rapid development of neural networks, data-driven grasping methods such as GPD[25] and PointNetGPD[13] adopt a two-stage framework. They first sample grasp candidates from point-cloud geometry and then classify these candidates using deep neural networks based on the local gripper closing region. However, due to geometric incompleteness in sparse point clouds, the sampled grasp candidates are often unreliable, and relying solely on the gripper closing region may further limit the accuracy and robustness of the subsequent evaluation module.
@@
To address these challenges, we propose SuperGrasp, a two-stage grasping pipeline similar to PointNetGPD[13], but with novel modules in both stages. Specifically, in the first stage, we directly estimate the superquadric coefficients from the object’s single-view point cloud to approximate its overall geometry. We then perform global similarity matching against an offline primitive database built on superquadric coefficients, where both shape and scale are considered, to enhance its geometric representation. The grasp poses of the top-$N$ most similar primitive objects are subsequently transformed into the point cloud of the target object, producing rich and effective grasp candidates.
In the second stage, to enable more reliable evaluation and refinement, we design an anchor-guided evaluation-refinement network based on PointNet++[23] as the backbone. The network enlarges the grasp-aware region and takes the initial grasp closure local region as an anchor region. Based on this design, it jointly analyzes the geometric features within the initial grasp closure region and the surrounding spatial geometric context, enabling more accurate and robust prediction of grasp feasibility, as well as refinement feasibility.
To validate the effectiveness of the proposed method, we conduct extensive grasping experiments in both simulation and real-world environments across multiple scenarios. The experimental results demonstrate that our method achieves strong stability while maintaining good generalization capability. The task pipeline is shown in Fig. 1.
The main contributions of this work are as follows:
  • We propose a superquadric-based similarity matching method that approximates the overall geometry of a target object from its incomplete point cloud using superquadric coefficients, retrieves multiple globally similar primitives from a database to enhance its geometric representation, and transfers grasp candidates from the matched primitives to the target object.
  • We propose an anchor-guided evaluation-refinement network that builds a grasp-centered contextual representation by integrating the initial closing region with its surrounding spatial neighborhood, enabling more reliable grasp assessment and local refinement.
@@ II Related Work
Grasping has long been a focus of robotics research. With the rapid development of computer vision and neural networks, grasping methods have become increasingly diverse. In this section, we introduce superquadrics and discuss two primary research directions in robotic grasping[19].
@@ II-A Superquadrics
To fully characterize a superquadric in 3D space, 5 geometric parameters and a 6-DoF pose are required:
$\{a_{x},a_{y},a_{z},\epsilon_{1},\epsilon_{2},\bm{T_{sq}}\},$
where $\bm{T_{sq}}=(\bm{R_{sq}},\bm{t_{sq}})$ denotes the pose of the superquadric, $R_{sq}$} represents its orientation, and $t_{sq}$} denotes its translation. Superquadrics are widely applied in object modeling, shape approximation, and scene understanding due to their low-dimensional representation and strong geometric interpretability. They are particularly effective for modeling objects with nearly convex structures, and the recovered parameters can directly convey the graphical, orientational, and shape characteristics of the objects.
@@ II-B Two-Stage Grasping
Two-stage grasping methods typically decompose the problem into grasp candidate generation and grasp evaluation. GPD[25] first samples grasp candidates in 3D space, labels them using force closure analysis [20], and trains a CNN to predict grasp feasibility from grasp-aligned projections. Method [13] further improves this paradigm by directly encoding the point cloud within the grasp region using a PointNet-based architecture for grasp quality prediction. Later, 3DSGrasp[17] introduced 3D reconstruction to compensate for missing geometric information in point clouds, thus improving both the generation of grasp candidates and the accuracy of grasp feasibility analysis. However, due to the continuous and high-dimensional nature of the grasp space, exhaustive spatial sampling is inherently time-consuming and inefficient. To address this issue, methods[3, 7] proposed to generate grasp candidates through similarity matching by directly matching the observed partial point cloud to database objects. However, operating directly on incomplete observations makes the overall matching and alignment processes less robust, thereby limiting the reliability of the generated grasp candidates.
@@ II-C End-to-End Grasping
To address sampling inefficiency, methods [6, 26, 12, 24, 15, 21] proposed end-to-end network frameworks that take point clouds as input and directly output grasp candidates with quality scores, relying on large-scale training data to ensure good generalization. Methods [27, 30] further introduced grasp refinement strategies, first analyzing global point cloud information to obtain coarse grasp poses, then refining them using local geometric details, fully leveraging point cloud spatial information and improving grasp success rates. Method [18] proposed VAE-based models that encode point clouds into implicit features and decode them to generate 6-DoF grasp poses. However, the aforementioned methods still exhibit limited generalization capabilities when handling variations in scene and object shapes.
@@ III Problem Formulation
We formulate the task as tabletop decluttering with a robotic arm equipped with a parallel-jaw gripper, although the proposed framework can also be applied to target-object grasping. Notably, only a single fixed depth camera is deployed within a $50\,\mathrm{cm}\times 50\,\mathrm{cm}$ workspace to capture the scene point cloud. The model takes as input the object point cloud together with the global scene point cloud $X\in\mathbb{R}^{n\times 3}$, and outputs a grasp pose $g$, defined as:
@@ IV-A Overview
To improve the efficiency and reliability of grasping, we propose SuperGrasp, a new two-stage framework. First, we introduce a geometric similarity matching strategy to generate grasp candidates (see Sec. IV-B). Second, we propose an end-to-end evaluation and refinement network, termed E-RNet, to evaluate the stability of the grasp and refine it locally (see Sec. IV-D).
@@ IV-B Superquadric Similarity Matching
Superquadrics provide an overall parametric representation of an object using only five shape parameters, $\{a_{x},a_{y},a_{z},\epsilon_{1},\epsilon_{2}\}$, and a 6-DoF pose, $\bm{T}_{sq}$, thereby enabling an alternative grasping paradigm. However, in the single-view setting, point clouds are often incomplete, making it difficult to accurately estimate superquadric coefficients. This leads to biases in both shape and scale, resulting in unstable grasp candidate quality and insufficient coverage. To address this issue, we propose a superquadric-based similarity matching method, as illustrated in Fig.4.
@@
We first build a database of 1.2k standard geometric primitives (e.g., cuboids, cylinders, frustums, and elliptical cylinders), as shown in Fig. 3. To better cover object scales, we apply proportional scaling to all primitives. For each object, we precompute a set of pre-grasp poses $G_{1}^{n}=\{g_{1},g_{2},\ldots\}$ using family-specific geometric templates that cover the main feasible grasping directions and positions. For each configuration, the corresponding gripper width $w$ is obtained by closing the gripper from an initially open state until the minimum distance between either inner finger surface and the object surface reaches a fixed threshold $\delta=0.0125\,\mathrm{m}$. Depending on the primitive geometry and scale, each object contributes 20 to 400 pre-grasp poses, with an average of about 200. We further precompute the superquadric coefficients of all database objects using the EMS (Expectation-Maximization Search) fitting method [14] and store them offline for subsequent retrieval.
Figure 3: The primitive object database constructed in this work, which contains 1.2k standard geometric objects of various shapes and scales, including cylinders, frustums, elliptical cylinders, and cuboids.
@@
Given a target object, we likewise use EMS to fit its superquadric and obtain the scale parameters $(a_{x},a_{y},a_{z})$ and shape parameters $(\epsilon_{1},\epsilon_{2})$. To separately quantify overall size and aspect ratio, we calculate the geometric mean of the scale parameters as follows:
$m=(a_{x}a_{y}a_{z})^{1/3},$ (2)
@@
$r=\left[\frac{a_{x}}{m},\frac{a_{y}}{m},\frac{a_{z}}{m}\right].$ (3)
@@
Figure 4: Overview: Our framework consists of two stages. In the first stage, we first estimate the superquadric coefficients of the target object and compute their similarity to those of the objects in the database. We then select the top-$N$ most similar objects and transfer their grasp poses to the target object through a transformation matrix $T$, followed by a coarse filtering step to remove noisy candidates. In the second stage, we enlarge the grasp-aware region of each initial grasp and feed the cropped regional point cloud into a PointNet++ network for feature extraction. We then further crop the features corresponding to the points within the initial gripper closing region and aggregate them via max pooling to obtain the grasp anchor feature. Finally, the anchor feature is fed into an evaluation network and a refinement network to predict grasp feasibility and refinement feasibility, respectively.
@@
$d_{\mathrm{scale}}=\left|\log\frac{m^{(q)}}{m^{(c)}}\right|,$ (7)
@@
We rank the remaining candidates in ascending order of $s$ and select the top-$K_{2}$ most similar objects. Their pre-grasp poses are then transferred to the target-object frame using the relative rigid transform $T=T_{sq}^{q}(T_{sq}^{c})^{-1}$}, where $T_{sq}^{q}$ and $T_{sq}^{c}$ denote the fitted poses of the query and candidate superquadrics, respectively. In all experiments, we set $w_{r}=1.0$, $w_{\epsilon}=1.2$, $\lambda_{s}=1.6$, $\lambda_{a}=2.0$, $K_{1}=50$ and $K_{2}=5$.
@@
We perform an additional coarse filtering step to remove unreasonable candidates, including: i) those for which the gripper collides with the target object body, ii) those that contain fewer than 50 target points between the gripper fingers, and iii) those for which the angle between the gripper closing direction and the surface normal at the fingertip contact points exceeds $20^{\circ}$.
This module yields rich and reliable grasp candidates for the subsequent evaluation and refinement module.
@@ IV-C Data Collection
For label assignment, unlike methods[13, 25], which grade grasp poses based on analytic metrics, we directly determine the label according to the execution outcome in simulation. Specifically, during execution, if the gripper collides with surrounding objects, the label is directly set to 0. If the gripper successfully lifts the target object and then reorients it to a pose with the gripper approach axis aligned vertically downward, and the grasp remains stable under a downward disturbance force of $1\,\mathrm{N}$ applied for $0.5\,\mathrm{s}$, as shown in Fig. 5, the label is set to 1. All other cases are labeled as 0.
@@ IV-D E-RNet: Evaluation and Refinement Network
between the gripper fingers and the local surface normals of the object, a certain degree of relative motion may occur between the gripper and the object during closure. Therefore, it is insufficient to consider only the grasp closing region; the spatial information around this region is also important for reliable grasp evaluation. In addition, although the first stage provides abundant grasp candidates, dense clutter and geometric incompleteness may still cause all initial candidates to fail, in which case slight pose adjustment can improve grasping adaptability. To address these issues, we propose a novel evaluation and refinement network, termed E-RNet. The overall architecture is illustrated in Fig. 4.
To enrich grasp-aware perception, we enlarge the initial grasp closing region. Specifically, the initial grasp pose is first deepened by $0.008\,\mathrm{m}$ along the approach direction, and then additionally rotated around the approach axis by $15^{\circ}$ and $15^{\circ}$. For stable learning, we transform the global point cloud into the gripper coordinate frame and use the enlarged closing region to crop the region of interest from the global point cloud, yielding a cropped point cloud of size $N\times 3$. The cropped point cloud is then fed into a PointNet++ backbone to extract point-wise features of size $N\times 128$.
@@
Unlike methods that fine-tune all grasp candidates in a multi-dimensional pose space to further expand the candidate set[11], we instead take the initial grasp region as a spatial anchor and explore the grasp feasibility of its neighboring regions, which is more efficient and time-saving. Moreover, unlike Regnet[30], which predicts residual corrections with respect to its matched single ground-truth grasp pose, our training strategy allows multiple valid refinement choices, as long as they can lead to a stable grasp. This design provides more refinement flexibility during inference.
@@
$\mathcal{L}_{\mathrm{r}}=$ $-\frac{1}{3M}\sum_{i=1}^{M}\sum_{j=1}^{3}\Big[y_{ij}\log\sigma(z_{ij})$ (11)